OpenStudy (anonymous):

3. What is the slope of the line? y=5x+2 A. 2 B. -2 C. 5 D. -5 4. What is the y-intercept of the line? y=5x+2 A. 2 B. -2 C. 5 D. -5 5. What is the first step to graphing a line using Slope-Intercept form? A. Plot the y-intercept B. Find a second point by counting the slope as rise over run. C. Draw a line between the two points

jimthompson5910 (jim_thompson5910):

In general, the slope of y = mx+b is m. The y-intercept is b. Example: y = 4x + 20 slope: 4 or 4/1 y-intercept: 20, which is the point (0, 20)

jimthompson5910 (jim_thompson5910):

this is because in this example, m = 4 and b = 20

OpenStudy (anonymous):

sorry im a little lost

jimthompson5910 (jim_thompson5910):

does my example make sense?

jimthompson5910 (jim_thompson5910):

in y = 4x + 20, compare that to y = mx+b to see that m = 4 and b = 20 so slope is 4 or 4/1 y-intercept is 20, which is the point (0, 20)

OpenStudy (anonymous):

not really sorry this is all confusing to me

jimthompson5910 (jim_thompson5910):

where are you stuck?

OpenStudy (anonymous):

all.

jimthompson5910 (jim_thompson5910):

have a look at this page and tell me if it helps or not http://www.purplemath.com/modules/strtlneq.htm